A researcher studying the nutrional value of a new candy places a 3.40 g sample of the candy inside a bomb calorimeter and combusts it in excess oxygen. The observed temperature increases is 2.17 C. If the heat capacity of the calorimeter is 30.40 kJ•K-1, how many nutrional Calories are there per gram of the candy?

Calories per gram of candy: ________ Cal/g
Answer #1

The increase in temperature can be written as \DeltaT = 2.17 oC or 2.17 K

Heat capacity of the calorimeter (C) = 30.4 kJ/K = (30.4/4.184) kcal/K = 7.26577 kcal/K = 7265.77 cal/K

The calories of candy = C * \DeltaT = 7265.77 cal/K * 2.17 K = 15766.73 cal

Therefore, the calories per gram of candy = 15766.73 cal/3.4 g = 4637.3 cal/g